5. Oktober 2007 — Freitag

Admin

  • WebCT: monitor system from 3:40 to 5:30 to see if yesterday’s changes fixed the problem
  • LDAP:
    • JOIN failures fixed
    • Feed changes requested from SOR owners
  • Backup:
    • DataDomain 31715: reboot 460 to ensure it is correctly configured
    • EMC 20242686: download replacement nsrmmd binary
    • Meeting with COMIS about backups for them

Comments are closed.